Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Geoff Cooney Wilkinstown Navan Co Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Formal Protest  23rd April 2002

A Chara,

Patrick O’Brien, Fri, 19/04/02 @ 20:59

Hi there. I wish to inform you that due to some requests, it has been decided to run a 3 x 40 event on Saturday 27th April, Start 11:00am. This event will be held in Fassaroe range. On Sunday there will be a prone event, first detail 10:00am Last entries 11:00am. This event will be held in the Dublin Rifle Club range. (Kilbride, Blessington, Co. Wicklow) These events will be registered. These events are to allow a person an additional opportunity of setting a qualifying score for selection purposes. There will be no prizes, but there will be an entry fee to cover costs.

Pat, Given these dates and venues, I have to submit a formal complaint as DURC Secretary, as follows:

DURC officially protests

1. the holding of any registered shoot on the same day as the Intervarsity,

2. the selection of DRC as the venue for the registered shoot on Sunday, April 28th.

On point 1:

The official Irish Intervarsity event, as recognized by CUSAI, is to be held in WTSC on Saturday 27th. This event has traditionally been a TCD/UCD event with a heavy social emphasis - but its recognition by CUSAI makes it the official Irish Intervarsity shooting event, and therefore the most important collegiate event of the year. At present, with only two college clubs, this may seem a trivial difference from the old Colours match – however it is more subtle in that prior to this year only UCD and TCD could compete - UL, DCU, DIT or any other third-level institution was not permitted to do so. Now, however, any third-level institute may enter the Intervarsity to compete. This has serious implications for college clubs, including financial ones in that sporting performance does affect capitation grants, the primary source of funding for college clubs. Since such clubs are the largest in the NTSA, a cynical attitude to the importance of this event is simply unacceptable.

We have had a squad of 16 people training regularly in both the gym and the range in preparation for this event. In addition, several of these shooters have been in training for selection to go to Pilsen for the first World Student Shooting Games. Pressures on time and venue space are self-evident and we do not hold the NTSA solely responsible for them: but to schedule a registered shoot on a day when another important competition is scheduled (even after that event was rescheduled to cater to the requirements of the national squad) is not acceptable to us.

On point 2:

We do not feel the DRC range is currently of the required standard to hold a registered shoot. It is in contravention of several ISSF regulations pertaining to range standards :

Rule number Violation

6.3.5.1.2 DRC is constructed facing south and thus the sun is almost always going to be in the shooter’s line of sight

6.3.7 DRC has no wind flags. In particular, note 6.3.7.2 and 7.4.1.2, which prohibits private wind flags; and 6.3.7.3 which requires a range to have two windflags, one at 10m downrange and one at 40m downrange, to be placed between each and every firing point. Further, wind baffles should be installed in accordance with 6.3.12.2.7 due to the severity of wind conditions which have been experienced in DRC in the past.

3.5.2 While no Irish range is at the standards of ISSF World Cup requirements, DRC meets only one requirement of the 15 stipulated within 3.5.2 - that of shelter for shooters and officials. 3.5.2.15 can be argued to be filled, but only cynically as the only parking facilities are the ditches on the access road. While we do not feel this is sufficient grounds for disallowing registered shoots in DRC, we do wish to express our desire to see DRC fulfill 3.5.2 as far as is reasonably possible.

While 6.3.5.1.1 allows technical committee members to allow small deviations, we feel that the above fall outside this description and should disallow DRC as a site for registered events at the present time, as the standard of the range is prejudicial against shooters attempting to shoot qualifying scores at DRC. While it is currently outside the resources of any club in the Republic to construct an international-standard range, that does not mean that international-standard firing lanes should be used for registered shoots.

We do not believe that this is in any way the fault of any member of DRC : nor do we believe these problems are insurmountable. DRC’s range is a valuable resource to both DRC and the target shooting community in general, and we welcome it’s addition to the list of Irish smallbore ranges. DRC are to be congratulated on successfully relocating after their deplorable experiences with their previous range and the negotiations for alternate sites. We wish to explicitly state that this protest is not to be taken as a personal criticism of any member of DRC or the NTSA.

range in Kilbride and look forward to shooting there once it has been completed, we officially protest the holding of the proposed registered shoot there on April 28th, and request that an alternate venue and/or date be selected.

However, we do not feel that shooters should suffer from the disadvantages of shooting in DRC’s range in registered shoots in its present state. Therefore, while we welcome the opening of the DRC Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

P J Harding 378 Ryevale Lawns Leixlip Co Kildare

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Formal Protest  23rd April 2002

A Chara,

Patrick O’Brien, Fri, 19/04/02 @ 20:59

Hi there. I wish to inform you that due to some requests, it has been decided to run a 3 x 40 event on Saturday 27th April, Start 11:00am. This event will be held in Fassaroe range. On Sunday there will be a prone event, first detail 10:00am Last entries 11:00am. This event will be held in the Dublin Rifle Club range. (Kilbride, Blessington, Co. Wicklow) These events will be registered. These events are to allow a person an additional opportunity of setting a qualifying score for selection purposes. There will be no prizes, but there will be an entry fee to cover costs.

Pat, Given these dates and venues, I have to submit a formal complaint as DURC Secretary, as follows:

DURC officially protests

1. the holding of any registered shoot on the same day as the Intervarsity,

2. the selection of DRC as the venue for the registered shoot on Sunday, April 28th.

On point 1:

The official Irish Intervarsity event, as recognized by CUSAI, is to be held in WTSC on Saturday 27th. This event has traditionally been a TCD/UCD event with a heavy social emphasis - but its recognition by CUSAI makes it the official Irish Intervarsity shooting event, and therefore the most important collegiate event of the year. At present, with only two college clubs, this may seem a trivial difference from the old Colours match – however it is more subtle in that prior to this year only UCD and TCD could compete - UL, DCU, DIT or any other third-level institution was not permitted to do so. Now, however, any third-level institute may enter the Intervarsity to compete. This has serious implications for college clubs, including financial ones in that sporting performance does affect capitation grants, the primary source of funding for college clubs. Since such clubs are the largest in the NTSA, a cynical attitude to the importance of this event is simply unacceptable.

We have had a squad of 16 people training regularly in both the gym and the range in preparation for this event. In addition, several of these shooters have been in training for selection to go to Pilsen for the first World Student Shooting Games. Pressures on time and venue space are self-evident and we do not hold the NTSA solely responsible for them: but to schedule a registered shoot on a day when another important competition is scheduled (even after that event was rescheduled to cater to the requirements of the national squad) is not acceptable to us.

On point 2:

We do not feel the DRC range is currently of the required standard to hold a registered shoot. It is in contravention of several ISSF regulations pertaining to range standards :

Rule number Violation

6.3.5.1.2 DRC is constructed facing south and thus the sun is almost always going to be in the shooter’s line of sight

6.3.7 DRC has no wind flags. In particular, note 6.3.7.2 and 7.4.1.2, which prohibits private wind flags; and 6.3.7.3 which requires a range to have two windflags, one at 10m downrange and one at 40m downrange, to be placed between each and every firing point. Further, wind baffles should be installed in accordance with 6.3.12.2.7 due to the severity of wind conditions which have been experienced in DRC in the past.

3.5.2 While no Irish range is at the standards of ISSF World Cup requirements, DRC meets only one requirement of the 15 stipulated within 3.5.2 - that of shelter for shooters and officials. 3.5.2.15 can be argued to be filled, but only cynically as the only parking facilities are the ditches on the access road. While we do not feel this is sufficient grounds for disallowing registered shoots in DRC, we do wish to express our desire to see DRC fulfill 3.5.2 as far as is reasonably possible.

While 6.3.5.1.1 allows technical committee members to allow small deviations, we feel that the above fall outside this description and should disallow DRC as a site for registered events at the present time, as the standard of the range is prejudicial against shooters attempting to shoot qualifying scores at DRC. While it is currently outside the resources of any club in the Republic to construct an international-standard range, that does not mean that international-standard firing lanes should be used for registered shoots.

We do not believe that this is in any way the fault of any member of DRC : nor do we believe these problems are insurmountable. DRC’s range is a valuable resource to both DRC and the target shooting community in general, and we welcome it’s addition to the list of Irish smallbore ranges. DRC are to be congratulated on successfully relocating after their deplorable experiences with their previous range and the negotiations for alternate sites. We wish to explicitly state that this protest is not to be taken as a personal criticism of any member of DRC or the NTSA.

However, we do not feel that shooters should suffer from the disadvantages of shooting in DRC’s range in registered shoots in its present state. Therefore, while we welcome the opening of the DRC Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Pat and Denise O’Brien Golden Hill Manor Kilbride Blessington Co Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Formal Protest  23rd April 2002

A Chara,

Patrick O’Brien, Fri, 19/04/02 @ 20:59

Hi there. I wish to inform you that due to some requests, it has been decided to run a 3 x 40 event on Saturday 27th April, Start 11:00am. This event will be held in Fassaroe range. On Sunday there will be a prone event, first detail 10:00am Last entries 11:00am. This event will be held in the Dublin Rifle Club range. (Kilbride, Blessington, Co. Wicklow) These events will be registered. These events are to allow a person an additional opportunity of setting a qualifying score for selection purposes. There will be no prizes, but there will be an entry fee to cover costs.

Pat, Given these dates and venues, I have to submit a formal complaint as DURC Secretary, as follows:

DURC officially protests

1. the holding of any registered shoot on the same day as the Intervarsity,

2. the selection of DRC as the venue for the registered shoot on Sunday, April 28th.

On point 1:

The official Irish Intervarsity event, as recognized by CUSAI, is to be held in WTSC on Saturday 27th. This event has traditionally been a TCD/UCD event with a heavy social emphasis - but its recognition by CUSAI makes it the official Irish Intervarsity shooting event, and therefore the most important collegiate event of the year. At present, with only two college clubs, this may seem a trivial difference from the old Colours match – however it is more subtle in that prior to this year only UCD and TCD could compete - UL, DCU, DIT or any other third-level institution was not permitted to do so. Now, however, any third-level institute may enter the Intervarsity to compete. This has serious implications for college clubs, including financial ones in that sporting performance does affect capitation grants, the primary source of funding for college clubs. Since such clubs are the largest in the NTSA, a cynical attitude to the importance of this event is simply unacceptable.

We have had a squad of 16 people training regularly in both the gym and the range in preparation for this event. In addition, several of these shooters have been in training for selection to go to Pilsen for the first World Student Shooting Games. Pressures on time and venue space are self-evident and we do not hold the NTSA solely responsible for them: but to schedule a registered shoot on a day when another important competition is scheduled (even after that event was rescheduled to cater to the requirements of the national squad) is not acceptable to us.

On point 2:

We do not feel the DRC range is currently of the required standard to hold a registered shoot. It is in contravention of several ISSF regulations pertaining to range standards :

Rule number Violation

6.3.5.1.2 DRC is constructed facing south and thus the sun is almost always going to be in the shooter’s line of sight

6.3.7 DRC has no wind flags. In particular, note 6.3.7.2 and 7.4.1.2, which prohibits private wind flags; and 6.3.7.3 which requires a range to have two windflags, one at 10m downrange and one at 40m downrange, to be placed between each and every firing point. Further, wind baffles should be installed in accordance with 6.3.12.2.7 due to the severity of wind conditions which have been experienced in DRC in the past.

3.5.2 While no Irish range is at the standards of ISSF World Cup requirements, DRC meets only one requirement of the 15 stipulated within 3.5.2 - that of shelter for shooters and officials. 3.5.2.15 can be argued to be filled, but only cynically as the only parking facilities are the ditches on the access road. While we do not feel this is sufficient grounds for disallowing registered shoots in DRC, we do wish to express our desire to see DRC fulfill 3.5.2 as far as is reasonably possible.

While 6.3.5.1.1 allows technical committee members to allow small deviations, we feel that the above fall outside this description and should disallow DRC as a site for registered events at the present time, as the standard of the range is prejudicial against shooters attempting to shoot qualifying scores at DRC. While it is currently outside the resources of any club in the Republic to construct an international-standard range, that does not mean that international-standard firing lanes should be used for registered shoots.

We do not believe that this is in any way the fault of any member of DRC : nor do we believe these problems are insurmountable. DRC’s range is a valuable resource to both DRC and the target shooting community in general, and we welcome it’s addition to the list of Irish smallbore ranges. DRC are to be congratulated on successfully relocating after their deplorable experiences with their previous range and the negotiations for alternate sites. We wish to explicitly state that this protest is not to be taken as a personal criticism of any member of DRC or the NTSA.

However, we do not feel that shooters should suffer from the disadvantages of shooting in DRC’s range in registered shoots in its present state. Therefore, while we welcome the opening of the DRC Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Nick West 95 Parkmore Drive Terenure Dublin 6W

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Formal Protest  23rd April 2002

A Chara,

Patrick O’Brien, Fri, 19/04/02 @ 20:59

Hi there. I wish to inform you that due to some requests, it has been decided to run a 3 x 40 event on Saturday 27th April, Start 11:00am. This event will be held in Fassaroe range. On Sunday there will be a prone event, first detail 10:00am Last entries 11:00am. This event will be held in the Dublin Rifle Club range. (Kilbride, Blessington, Co. Wicklow) These events will be registered. These events are to allow a person an additional opportunity of setting a qualifying score for selection purposes. There will be no prizes, but there will be an entry fee to cover costs.

Pat, Given these dates and venues, I have to submit a formal complaint as DURC Secretary, as follows:

DURC officially protests

1. the holding of any registered shoot on the same day as the Intervarsity,

2. the selection of DRC as the venue for the registered shoot on Sunday, April 28th.

On point 1:

The official Irish Intervarsity event, as recognized by CUSAI, is to be held in WTSC on Saturday 27th. This event has traditionally been a TCD/UCD event with a heavy social emphasis - but its recognition by CUSAI makes it the official Irish Intervarsity shooting event, and therefore the most important collegiate event of the year. At present, with only two college clubs, this may seem a trivial difference from the old Colours match – however it is more subtle in that prior to this year only UCD and TCD could compete - UL, DCU, DIT or any other third-level institution was not permitted to do so. Now, however, any third-level institute may enter the Intervarsity to compete. This has serious implications for college clubs, including financial ones in that sporting performance does affect capitation grants, the primary source of funding for college clubs. Since such clubs are the largest in the NTSA, a cynical attitude to the importance of this event is simply unacceptable.

We have had a squad of 16 people training regularly in both the gym and the range in preparation for this event. In addition, several of these shooters have been in training for selection to go to Pilsen for the first World Student Shooting Games. Pressures on time and venue space are self-evident and we do not hold the NTSA solely responsible for them: but to schedule a registered shoot on a day when another important competition is scheduled (even after that event was rescheduled to cater to the requirements of the national squad) is not acceptable to us.

On point 2:

We do not feel the DRC range is currently of the required standard to hold a registered shoot. It is in contravention of several ISSF regulations pertaining to range standards :

Rule number Violation

6.3.5.1.2 DRC is constructed facing south and thus the sun is almost always going to be in the shooter’s line of sight

6.3.7 DRC has no wind flags. In particular, note 6.3.7.2 and 7.4.1.2, which prohibits private wind flags; and 6.3.7.3 which requires a range to have two windflags, one at 10m downrange and one at 40m downrange, to be placed between each and every firing point. Further, wind baffles should be installed in accordance with 6.3.12.2.7 due to the severity of wind conditions which have been experienced in DRC in the past.

3.5.2 While no Irish range is at the standards of ISSF World Cup requirements, DRC meets only one requirement of the 15 stipulated within 3.5.2 - that of shelter for shooters and officials. 3.5.2.15 can be argued to be filled, but only cynically as the only parking facilities are the ditches on the access road. While we do not feel this is sufficient grounds for disallowing registered shoots in DRC, we do wish to express our desire to see DRC fulfill 3.5.2 as far as is reasonably possible.

While 6.3.5.1.1 allows technical committee members to allow small deviations, we feel that the above fall outside this description and should disallow DRC as a site for registered events at the present time, as the standard of the range is prejudicial against shooters attempting to shoot qualifying scores at DRC. While it is currently outside the resources of any club in the Republic to construct an international-standard range, that does not mean that international-standard firing lanes should be used for registered shoots.

We do not believe that this is in any way the fault of any member of DRC : nor do we believe these problems are insurmountable. DRC’s range is a valuable resource to both DRC and the target shooting community in general, and we welcome it’s addition to the list of Irish smallbore ranges. DRC are to be congratulated on successfully relocating after their deplorable experiences with their previous range and the negotiations for alternate sites. We wish to explicitly state that this protest is not to be taken as a personal criticism of any member of DRC or the NTSA.

However, we do not feel that shooters should suffer from the disadvantages of shooting in DRC’s range in registered shoots in its present state. Therefore, while we welcome the opening of the DRC Misé le meas,

Mark Dennehy, DURC Secretary

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Sandy Wilson 9 Abbey Court Monkstown Co Dublin

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Formal Protest  23rd April 2002

A Chara,

Patrick O’Brien, Fri, 19/04/02 @ 20:59

Hi there. I wish to inform you that due to some requests, it has been decided to run a 3 x 40 event on Saturday 27th April, Start 11:00am. This event will be held in Fassaroe range. On Sunday there will be a prone event, first detail 10:00am Last entries 11:00am. This event will be held in the Dublin Rifle Club range. (Kilbride, Blessington, Co. Wicklow) These events will be registered. These events are to allow a person an additional opportunity of setting a qualifying score for selection purposes. There will be no prizes, but there will be an entry fee to cover costs.

Pat, Given these dates and venues, I have to submit a formal complaint as DURC Secretary, as follows:

DURC officially protests

1. the holding of any registered shoot on the same day as the Intervarsity,

2. the selection of DRC as the venue for the registered shoot on Sunday, April 28th.

On point 1:

The official Irish Intervarsity event, as recognized by CUSAI, is to be held in WTSC on Saturday 27th. This event has traditionally been a TCD/UCD event with a heavy social emphasis - but its recognition by CUSAI makes it the official Irish Intervarsity shooting event, and therefore the most important collegiate event of the year. At present, with only two college clubs, this may seem a trivial difference from the old Colours match – however it is more subtle in that prior to this year only UCD and TCD could compete - UL, DCU, DIT or any other third-level institution was not permitted to do so. Now, however, any third-level institute may enter the Intervarsity to compete. This has serious implications for college clubs, including financial ones in that sporting performance does affect capitation grants, the primary source of funding for college clubs. Since such clubs are the largest in the NTSA, a cynical attitude to the importance of this event is simply unacceptable.

We have had a squad of 16 people training regularly in both the gym and the range in preparation for this event. In addition, several of these shooters have been in training for selection to go to Pilsen for the first World Student Shooting Games. Pressures on time and venue space are self-evident and we do not hold the NTSA solely responsible for them: but to schedule a registered shoot on a day when another important competition is scheduled (even after that event was rescheduled to cater to the requirements of the national squad) is not acceptable to us.

On point 2:

We do not feel the DRC range is currently of the required standard to hold a registered shoot. It is in contravention of several ISSF regulations pertaining to range standards :

Rule number Violation

6.3.5.1.2 DRC is constructed facing south and thus the sun is almost always going to be in the shooter’s line of sight

6.3.7 DRC has no wind flags. In particular, note 6.3.7.2 and 7.4.1.2, which prohibits private wind flags; and 6.3.7.3 which requires a range to have two windflags, one at 10m downrange and one at 40m downrange, to be placed between each and every firing point. Further, wind baffles should be installed in accordance with 6.3.12.2.7 due to the severity of wind conditions which have been experienced in DRC in the past.

3.5.2 While no Irish range is at the standards of ISSF World Cup requirements, DRC meets only one requirement of the 15 stipulated within 3.5.2 - that of shelter for shooters and officials. 3.5.2.15 can be argued to be filled, but only cynically as the only parking facilities are the ditches on the access road. While we do not feel this is sufficient grounds for disallowing registered shoots in DRC, we do wish to express our desire to see DRC fulfill 3.5.2 as far as is reasonably possible.

While 6.3.5.1.1 allows technical committee members to allow small deviations, we feel that the above fall outside this description and should disallow DRC as a site for registered events at the present time, as the standard of the range is prejudicial against shooters attempting to shoot qualifying scores at DRC. While it is currently outside the resources of any club in the Republic to construct an international-standard range, that does not mean that international-standard firing lanes should be used for registered shoots.

We do not believe that this is in any way the fault of any member of DRC : nor do we believe these problems are insurmountable. DRC’s range is a valuable resource to both DRC and the target shooting community in general, and we welcome it’s addition to the list of Irish smallbore ranges. DRC are to be congratulated on successfully relocating after their deplorable experiences with their previous range and the negotiations for alternate sites. We wish to explicitly state that this protest is not to be taken as a personal criticism of any member of DRC or the NTSA.

However, we do not feel that shooters should suffer from the disadvantages of shooting in DRC’s range in registered shoots in its present state. Therefore, while we welcome the opening of the DRC Misé le meas,

Mark Dennehy, DURC Secretary